the hydrated chloride of divalent metl loses 14.73%(by weight) of its mass by dehydration. if one mole of hydrated salt contain 2 moles of hydration,calculate weight% of chlorine in hydrated salt..

Grade:11

1 Answers

Arun
3 years ago
The molecular weight of water H2O=2(1)+16=18.
When 1 mole of CuSO45H2O is heated, 5 moles of water are lost.
The molecular weight of CuSO45H2O=64+32+4(16)+5(18)=250 g/mol
Mass of water lost =5(18)=90 g
Percentage loss of mass of CuSO45H2O=25090×100=36 %.